TheSmartOne (thesmartone):

f(x)=-2x^3-7x^2-166x+85 Find the complex zeros of f. Repeat any zeros if their multiplicity is greater than 1.

TheSmartOne (thesmartone):

@satellite73

TheSmartOne (thesmartone):

@jim_thompson5910 @zepdrix

TheSmartOne (thesmartone):

@ganeshie8 @myininaya

OpenStudy (anonymous):

yikes a third degree polynomial

OpenStudy (anonymous):

i have no idea how you are supposed to do this, but apparently the one real zero is at \(\frac{1}{2}\) which means it factors as \[-(2x-1)(\text{something})\]

OpenStudy (anonymous):

the "something" will be a quadratic with two complex zeros they have to be conjugates so they are different, making the multiplicity of each zero as 1

OpenStudy (anonymous):

me, i would cheat because this is donkey work http://www.wolframalpha.com/input/?i=-2x^3-7x^2-166x%2B85

TheSmartOne (thesmartone):

I also have to write it in factored form...

OpenStudy (anonymous):

yeah it is a thankless task to find the zeros of a third degree polynomial

OpenStudy (anonymous):

since on zero is \(\frac{1}{2}\) you can factor it as \[-(2 x-1) (x^2+4 x+85)\]

OpenStudy (anonymous):

* one zero

OpenStudy (anonymous):

the other two complex ones you find either using the quadratic formula to solve \[x^2+4x+85=0\] or, what is easier because 4 is even, completing the square

TheSmartOne (thesmartone):

I have an example So for the equation -2x^3-27x^2-116x+65 it was factored down to (-2x+1)(x-(-7+4i))(x-(-7-4i))

OpenStudy (anonymous):

similarly you can factor this one once you find the complex zeros

TheSmartOne (thesmartone):

so\[x^2+4x+85=0\] \[x^2+4x+4=-85+4\] \[(x+2)^2=-81\] \[x+2=\pm9i\] \[x=-2\pm9i\]

TheSmartOne (thesmartone):

So\[x=\frac{ 1 }{ 2 },-2+9i, -2-9i\]

TheSmartOne (thesmartone):

\[-2x^3-7x^2-166x+85=(2x-1)(x-(-2+9i))(x-(-2-9i))\]

TheSmartOne (thesmartone):

@satellite73

TheSmartOne (thesmartone):

I am not sure if I wrote it correctly in factored form. Is it (2x-1)(x-(-2+9i))(x-(-2-9i))

zepdrix (zepdrix):

Hmm your roots look good! Factored? mmm yesss good job.

zepdrix (zepdrix):

Hard to read through all those brackets sometimes lol

TheSmartOne (thesmartone):

Well it is telling me it is wrong...

zepdrix (zepdrix):

Hmm :d

TheSmartOne (thesmartone):

wait satellite said -(2x-1) so I guess I missed the - sign?

TheSmartOne (thesmartone):

I got it correct?

TheSmartOne (thesmartone):

*.. not a ?

zepdrix (zepdrix):

The negative floating out front shouldn't affect your roots 0_o that's very strange....

TheSmartOne (thesmartone):

I got it correct it should have been a (-2x+1) not (2x-1)

zepdrix (zepdrix):

Those are the same :U But whatever lol. They're being silly I think.

TheSmartOne (thesmartone):

No actually I forgot to put the - sign so I just distributed it in to it...

zepdrix (zepdrix):

I mean....about half way through the problem, after you find your 1/2 root, you get something of this form:\[\Large\rm 0=(2x-1)(-x^2-4x-85)\]Pulling a negative out of the bracketed part:\[\Large\rm 0=-(2x-1)(x^2+4x+85)\]And then just divide both sides by -1. I don't see why they want you to hold on to that lonely negative sign >.< So weird!

TheSmartOne (thesmartone):

Thanks again both of you @satellite73 and @zepdrix